OpenStudy (anonymous):

logbase2(x^2 + 6x) = 4

OpenStudy (anonymous):

\[\log_{2} (x^2+6x)=4\]

OpenStudy (anonymous):

i need to solve for x, but i have no idea how

OpenStudy (anonymous):

\[\log_bc=a\Rightarrow c=b^a\]

OpenStudy (anonymous):

hold on i think i got it, give me a minute

OpenStudy (anonymous):

x^2 + 6x = 16 is that right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

:)